The average number of people in the United States, who quit smoking cigarettes weekly is 96. Assume the standard deviation is 16 people. If random samples of 36 smokers are selected, find the probability that the mean number of people who quit smoking is 90 and 100 people.
Answer by stanbon(75887)

The average number of people in the United States, who quit smoking cigarettes weekly is 96. Assume the standard deviation is 16 people. If random samples of 36 smokers are selected, find the probability that the mean number of people who quit smoking is 90 and 100 people.
--------------------
t(90) = (90-96)/[16/sqrt(36)] = -6/(16/6) = -36/16 = -2.25
---
t(100) = (100-96)/[16/sqrt(36)] = 4/(16/6) = 24/16 = 3/2
-----
P(90<= x <=100) = P(-2.25<= t <=3/2) = tcdf(-2.25,1.5,35) = 0.9133
============================
Cheers,
Stan H.